OpenStudy (anonymous):

Re-write this equation in the standard form. y - 6 = 1.3(x + 7) a. 1.3x -10y = 151 c. 13x - 10y = -151 b. 13x +10y = -151 d. 1.3x - 10y = -151

OpenStudy (anonymous):

List the integers 1 to 10 inclusive that are even and a multiple of 3. a. 2, 4, 6, 8 b. 6 c. 6, 12

OpenStudy (anonymous):

U can't figure out which is the correct ansewr in the first question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

for the second one We have the set of integers from [1,10] so it's 1,2,3,4,5,6,7,8,9,10 the even ones are 2,4,6,8,10 the ones that are dividable with 3 is:6

OpenStudy (anonymous):

Five chips numbered 1 through 5 are placed in a bag. A chip is drawn and replaced. Then a second chip is drawn. Find the probability that the first chip is even AND the second chip is odd. a.4/25 b.9/25 c.6/25 d.12/25

OpenStudy (anonymous):

Write the point-slope form of an equation for a line that passes through the point with the given slope. (–4, 3), slope = 1 a. y – 3 = 1(x + 4) c. y – 3 = 1(x – 4) b. y + 3 = 1(x + 4) d. y – 3 = –(x + 4)

OpenStudy (anonymous):

do you know the formula for the slope

OpenStudy (anonymous):

yes but i just need help right now like fast so just give me the answers if you can please

OpenStudy (anonymous):

@jovel

OpenStudy (anonymous):

i leave from this class in 10 mins so give me whatever answers you have :(